# Learning Low-Dimensional Embeddings for Black-Box Optimization

### v1

@@ 4.3 Linear Decoder
In order to provide an interpretation of the Meta-GLIS algorithm, we modified the architecture of the autoencoder by considering a linear decoder $\mathrm{D}_{\phi_{d}}^{\mathrm{lin}}(z)$ of the form $\hat{x}=Az+b$, with $A\in\mathbb{R}^{n_{x}\times n_{z}}$ and $b\in\mathbb{R}^{n_{x}}$. The structure of the encoder and the dimension of the latent variable ($n_{z}=3$) are not changed with respect to the case discussed in the previous paragraph.
From Figs.  4 and 6, we can see that Meta-GLIS with linear decoder still outperforms original GLIS, with only a slight performance deterioration with respect to Meta-GLIS with a nonlinear decoder.
In Fig. 7, we can visualize the magnitude of the trained weights of the matrix $A$, after normalizing its columns to have norm $1$. In this way, we can see which MPC hyperparameters are more sensitive to a variation of the latent variable $z$. From the figure, we can see that variation of $z$ mainly influences three MPC hyperparameters, namely: $N_{p}$ and the log-tolerances of the QP solver $\mathrm{QP}_{\epsilon_{abs}}$, $\mathrm{QP}_{\epsilon_{rel}}$.
The relevance of each MPC hyperparameter is reflected on the empirical distributions of the sampled points during the optimization and of the optimal ones. These distributions are visualized in Fig. 5. We can observe that, in the $30$ iterations of the algorithm, GLIS tends to sample on the bounds of the input domains (shaded blue areas). This behavior is mitigated in Meta-GLIS (shaded orange and purple areas), where the exploration is more concentrated in the region containing the optima (like in $Q_{\Delta u}$, $W_{v}$), or else is more uniform within the research space ($q_{y}$, $W_{w_{33}}$, $W_{w_{33}}$). Looking at the explored points when using the linear decoder, the exploration remains uniform in the search space for those hyperparameters that are more sensitive to the variation of the latent variable (namely, $N_{p}$, $\mathrm{QP}_{\epsilon_{abs}}$, $\mathrm{QP}_{\epsilon_{rel}}$).
@@ 5 Conclusions
We have proposed a computationally efficient method for reducing the dimension of the input space where black-box optimization is performed. The core idea of our approach is to learn a low-dimensional embedding of the decision variable space using an autoencoder trained on precomputed solutions of samples of optimization problems within a given class. These precomputed solutions can be generated, for example, via simulations with varying parameters and, if needed, through time-consuming computations. Moving from simulation to real, optimization is performed in the embedded space using experiment-driven techniques such as Bayesian Optimization or GLIS. This approach significantly reduces the number of required experiments compared to optimizing directly in the original higher-dimensional space, as we have shown through Rosenbrock benchmark and automated calibration of MPC hyperparameters.
Future work will focus on extending the methodology to experiment-driven preference-based optimization, where objective functions cannot be directly measured but are instead evaluated through qualitative comparisons between the outputs of different experiments.

### v2

@@ 5.3 Linear Decoder
In order to provide an interpretation of the Meta-GLIS algorithm, we modified the architecture of the autoencoder by considering a linear decoder $\mathrm{D}_{\phi_{d}}^{\mathrm{lin}}(z)$ of the form $\hat{x}=Az+b$, with $A\in\mathbb{R}^{n_{x}\times n_{z}}$ and $b\in\mathbb{R}^{n_{x}}$. The structure of the encoder and the dimension of the latent variable ($n_{z}=3$) are not changed with respect to the case discussed in the previous paragraph. From Figs. 6 and 7, we can see that Meta-GLIS with linear decoder still outperforms original GLIS, with only a slight performance deterioration with respect to Meta-GLIS with a nonlinear decoder.
In Fig. 8, we can visualize the magnitude of the trained weights of the matrix $A$, after normalizing its columns to have unit Euclidean norm. In this way, we can see which MPC hyperparameters are more sensitive to a variation of the latent variable $z$. From the figure, we can see that variation of $z$ mainly influences three MPC hyperparameters, namely: $N_{p}$ and the log-tolerances of the QP solver $\mathrm{QP}_{\epsilon_{abs}}$, $\mathrm{QP}_{\epsilon_{rel}}$.
The relevance of each MPC hyperparameter is reflected on the empirical distributions of the sampled points during the optimization and of the optimal ones. These distributions are visualized in Fig. 5. We can observe that, in the $30$ iterations of the algorithm, GLIS tends to explore the bounds of the input domains (shaded blue areas). This behavior is mitigated in Meta-GLIS (shaded orange and purple areas), where the exploration is more concentrated in the region containing the optima (like in $Q_{\Delta u}$, $W_{v}$), or otherwise more uniform within the search space ($q_{y}$, $W_{w_{33}}$, $W_{w_{33}}$). Looking at the explored points when using the linear decoder, the exploration remains uniform in the search space for those hyperparameters that are more sensitive to the variation of the latent variable (namely, $N_{p}$, $\mathrm{QP}_{\epsilon_{abs}}$, $\mathrm{QP}_{\epsilon_{rel}}$).
@@ 6 Conclusions
We presented a method for reducing the dimension of the decision variable space in black-box optimization, thereby decreasing the computational effort required for numerical solution. The core idea of our approach is to learn a low-dimensional embedding within the original decision variable space using an autoencoder trained on precomputed solutions of optimization problems sampled within a given class. These precomputed solutions can be generated, for instance, through simulations with varying parameters and, if necessary, via computationally intensive optimization procedures. Moving from simulation to real, optimization is performed in the embedded space using experiment-driven techniques such as Bayesian Optimization or GLIS. This approach significantly reduces the number of required experiments compared to optimizing directly in the original higher-dimensional space, as we have shown on the Rosenbrock function benchmark and automated calibration of MPC hyperparameters.
Future research will extend the methodology in several directions. On the methodological side, we aim to integrate the approach with preference-based optimization, where objective functions cannot be directly measured but are instead evaluated through qualitative comparisons. On the theoretical side, further work is needed to sharpen the generalization bounds, as well as to extend the analysis to broader meta-learning scenarios such as amortized optimization and system identification in reduced spaces, ultimately providing theoretical tools to assess and enhance the trustworthiness of reduced-space methods.
Appendix
In this appendix we provide the proof of Theorem 1. First, a set of useful result is provided.
Property 1 (i.i.d. structure of the performance gap)
Let $(\theta^{(i)},\xi^{(i)},\omega^{(i)})$, $i=1,\ldots,m$, be i.i.d. draws from the product $p(\theta)\otimes p(\xi)\otimes p(\omega)$. Define the performance gaps
$\Psi^{(i)}\;:=\;\Psi\big(\theta^{(i)},\xi^{(i)},\omega^{(i)}\big),\qquad i=1,\dots,m,$ (23)
as in (9). Then $\{\Psi^{(i)}\}_{i=1}^{m}$ are i.i.d. real-valued random variables.
Proof of Property 1
By hypothesis, $(\theta^{(i)},\xi^{(i)},\omega^{(i)})$ are i.i.d. samples drawn from the product distribution $p(\theta)\otimes p(\xi)\otimes p(\omega)$. For each triple $(\theta,\xi,\omega)$, the performance gap $\Psi(\theta,\xi,\omega)$ is obtained by applying the same deterministic transformation, consisting of the optimization algorithms $\mathcal{A}_{x}$ and $\mathcal{A}_{z}$, the decoder $\mathrm{D}$, and the objective function $f$. Since applying a fixed deterministic transformation to i.i.d. inputs preserves both independence and identical distribution, the resulting values
$\Psi^{(i)}=\Psi(\theta^{(i)},\xi^{(i)},\omega^{(i)}),\qquad i=1,\dots,m,$ (24)
are themselves i.i.d. real-valued random variables. $\blacksquare$
The following lemma by Dvoretzky–Kiefer–Wolfowitz (DKW), and refined by Massart in [10] establishes that the empirical cumulative distribution function $\hat{F}_{m}$ of $m$ i.i.d. samples uniformly approximates the true cumulative distribution function $F$ of the underlying random variable $\Psi$. Specifically, with probability at least $1-\delta$, the maximum deviation $\sup_{t}|\hat{F}_{m}(t)-F(t)|$ is bounded by $\varepsilon_{m}=\sqrt{\log(2/\delta)/(2m)}$. This result provides a non-asymptotic, distribution-free guarantee on the accuracy of the empirical CDF.
Lemma 1 (DKW–Massart inequality [10])
Let $\Psi$ be a real-valued random variable with cumulative density function (CDF) $F$ and let $\hat{F}_{m}$ be the empirical CDF of i.i.d. samples $\Psi^{(1)},\ldots,\Psi^{(m)}$ defined as:
$\hat{F}_{m}(t)=\frac{1}{m}\sum_{i=1}^{m}\mathbf{1}[\Psi^{(i)}\leq t],\ \ t\in\mathbb{R},$ (25)
with $\mathbf{1}$ denoting the indicator function.
Then, for any $\delta\in(0,1)$, with
$\varepsilon_{m}\ :=\ \sqrt{\frac{\log(2/\delta)}{2m}},$ (26)
it holds that
$\mathbb{P}_{S}\!\left(\sup_{t\in\mathbb{R}}\,\big|\hat{F}_{m}(t)-F(t)\big|\ \leq\ \varepsilon_{m}\right)\ \geq\ 1-\delta,$ (27)
where $S:=\{\Psi^{(i)}\}_{i=1}^{m}$ is the set of i.i.d. samples of $\Psi$.
In the following section, we provide the proof of Theorem 1.
Proof of Theorem 1
We proceed in two steps.
Step 1. The goal of this step is to show that the true quantile $Q_{1-\alpha}:=\inf\{t:F(t)\geq 1-\alpha\}$ (for any $\alpha\in(0,1)$) can be upper bounded, with high probability (i.e., with probability at least $1-\delta$, for arbitrary $\delta\in(0,1)$), by a suitable order statistic of the validation samples $S=\{\Psi^{(i)}:=\Psi(\theta^{(i)},\xi^{(i)},\omega^{(i)})\}_{i=1}^{m}$, with
$(\theta^{(i)},\xi^{(i)},\omega^{(i)})\sim p(\theta)\otimes p(\xi)\otimes p(\omega).$ (28)
Let $t^{\star}:=\Psi_{(k^{\star})}$ be the $k^{\star}$-th order statistic of the validation set. By construction, at least $k^{\star}$ samples satisfy $\Psi^{(i)}\leq t^{\star}$, hence by definition of the empirical CDF in eq. (25), we have:
$\hat{F}_{m}(t^{\star})\ \geq\ \frac{k^{\star}}{m},$ (29)
where the inequality (and not equality) holds because multiple samples $\Psi^{(i)}$ may coincide with $t^{\star}$.
Define the event
$\mathcal{E}:=\Big\{\sup_{t\in\mathbb{R}}|\hat{F}_{m}(t)-F(t)|\leq\varepsilon_{m}\Big\}.$ (30)
On $\mathcal{E}$, we have
$F(t^{\star})\ \geq\ \hat{F}_{m}(t^{\star})-\varepsilon_{m}\ \geq\ \frac{k^{\star}}{m}-\varepsilon_{m}\ \geq\ 1-\alpha,$ (31)
where the last inequality follows from the definition of $k^{\star}$.
By definition of the true quantile, inequality (31) implies
$Q_{1-\alpha}\ \leq\ t^{\star}\ =\ \Psi_{(k^{\star})}.$ (32)
Let us set $\varepsilon_{m}:=\sqrt{\frac{\log(2/\delta)}{2m}}$. Since $\{\Psi^{(i)}\}_{i=1}^{m}$ are i.i.d. random variables (Property 1), the DKW–Massart inequality (Lemma 1) can be applied. This implies that for any $\delta\in(0,1)$, the event $\mathcal{E}$ holds with probability at least $1-\delta$ over the draw of the validation set $S$. Thus,
$\mathbb{P}_{S}\!\left(Q_{1-\alpha}\ \leq\ \Psi_{(k^{\star})}\right)\ \geq\ 1-\delta.$ (33)
Step 2. We now leverage the quantile bound from Step 1 to control the performance on a new draw $(\theta^{\rm new},\xi^{\rm new},\omega^{\rm new})\sim p(\theta)\otimes p(\xi)\otimes p(\omega)$. By definition of the $(1-\alpha)$-quantile,
$\mathbb{P}_{\Psi}\!\big(\Psi^{\rm new}\leq Q_{1-\alpha}\big)=1-\alpha.$ (34)
If $\mathcal{E}$ holds, then $Q_{1-\alpha}\leq\Psi_{(k^{\star})}$, and hence
$\{\Psi^{\rm new}\leq\Psi_{(k^{\star})}\}\ \supseteq\ \{\Psi^{\rm new}\leq Q_{1-\alpha}\}.$ (35)
Therefore,
$\mathbb{P}_{\Psi}\!\left(\Psi^{\rm new}\leq\Psi_{(k^{\star})}\,\Big|\,S\right)$ $\;\geq\;\mathbb{P}_{\Psi}\!\big(\Psi^{\rm new}\leq Q_{1-\alpha}\big)$
= $=\;1-\alpha.$ (36)
Taking the outer probability over the validation samples $S$, we obtain
$\mathbb{P}_{S}\!\Big(\mathbb{P}_{\Psi}\!\left(\Psi^{\rm new}\leq\Psi_{(k^{\star})}\,\Big|\,S\right)\geq 1-\alpha\Big)\;\geq\;1-\delta,$ (37)
which completes the proof. $\blacksquare$